# Source Prompt Disentangled Inversion for Boosting Image Editability with Diffusion Models

### v1

@@ 3.1 Analysis on DDIM Inversion
$z_{t}=C_{t,1}*z_{t-1}+C_{t,2}*{\color[rgb]{1,0,0}\epsilon_{\theta}(z_{t},t,c)},$ (1)
@@
The coupling of DDIM inversion with source prompt is rooted in the use of Eq. 2. In the ideal inversion (i.e., Eq. 1), the input should be $z_{t}$ and $t$, whereas previous methods have utilized $z_{t-1}$ and $t-1$ as the input (i.e., Eq. 2). However, $z_{t-1}$ is obtained by denoising $z_{t}$ conditioned with source prompt. Therefore, Eq. 2 actually introduces source prompt prior into the update. Taking the last inversion step as an example, the neural network should receive a latent $z_{T}$ a pure noise without any prior information as the input to obtain the ideal inversion code as $z_{T}=C_{T,1}*z_{T-1}+C_{T,2}*{\color[rgb]{1,0,0}\epsilon_{\theta}(z_{T},T,c)}$, while most previous methods utilize the latent feature $z_{T-1}$, which is obtained through one-step denoising conditioned on the source prompt, as the input, i.e., $\hat{z}_{T}=C_{T,1}*z_{T-1}+C_{T,2}*{\color[rgb]{0,0,1}\epsilon_{\theta}(z_{T-1},{T-1},c)}$. The coupling of the inverted noise code and source prompt is not limited to the final inversion step, as the multi-step generation nature of LDM accumulates the divergence between the inverted noise and the ideal noise. This eventually results in the inclusion of source prompt prior in the inverted noise, causing a notable deviation from the ideal noise, i.e., gap of noise $D_{Noi}$ as illustrated in Fig. 2(a). This deviation ultimately affects the editing stability and fidelity of the given real image.
@@ 3.2 Source Prompt Disentangled Inversion (SPDInv)
$x=f_{\theta}(x)\quad where\quad x=z_{t},f_{\theta}(x)=C_{t,2}*\epsilon_{\theta}(z_{t},t,c)+C_{t,1}*z_{t-1}.$ (3)
@@ Appendix 0.A Reduction of Noise Gap by SPDInv
We conduct experiments to evaluate the noise gap reduction by the proposed SPDInv (please refer to Fig. 2 in the main paper and the analysis in the introduction section for details of noise gap $D_{noi}$). We first utilize the captions extracted from the COCO2017 evaluation dataset[26] to generate 100 images by Stable Diffusion V1.4. We then record the initial noises $z_{T}$ and all the latent features $z_{t},t\in[0,T)$, which can be regarded as the ground-truth features of the generated images (please refer to the generation path in Fig. 2 of our main paper). Since previous text-driven image editing methods mostly employ DDIM inversion to compute the inverted noise code, we thus compare DDIM inversion with our SPDInv on their inversion performance using the generated ground-truth features.
During the inversion process, all inverted latent features are recorded. Then for each inversion step $t$, we can calculate the mean square error (MSE) between the ground-truth and inverted latent features. Finally, the MSEs of all the 100 images are averaged as the $D_{noi}$, which is illustrated in Fig. 7. Compared with DDIM inversion, our SPDInv reduces the noise gap from 0.06 to 0.04 ($36\%\downarrow$) after 50 inversion steps. Owe to the reduced noise gap, our SPDInv can effectively reduce the effect of source prompt on the inverted noise code, and consequently reduce the editing artifacts.
@@
$\begin{split}z_{t-1}&=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}z_{t}+\sqrt{\alpha_{t-1}}\bigg{(}\sqrt{\frac{1}{\alpha_{t-1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1}\bigg{)}\epsilon_{\theta}(z_{t},t,c),\\ \frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}z_{t}&=z_{t-1}-\sqrt{\alpha_{t-1}}\bigg{(}\sqrt{\frac{1}{\alpha_{t-1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1}\bigg{)}\epsilon_{\theta}(z_{t},t,c),\\ \frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}z_{t}&=z_{t-1}+\sqrt{\alpha_{t-1}}\bigg{(}\sqrt{\frac{1}{\alpha_{t}}-1}-\sqrt{\frac{1}{\alpha_{t-1}}-1}\bigg{)}\epsilon_{\theta}(z_{t},t,c),\\ z_{t}&=\frac{\sqrt{\alpha_{t}}}{\sqrt{\alpha_{t-1}}}z_{t-1}+\sqrt{\alpha_{t}}\bigg{(}\sqrt{\frac{1}{\alpha_{t}}-1}-\sqrt{\frac{1}{\alpha_{t-1}}-1}\bigg{)}{\color[rgb]{1,0,0}\epsilon_{\theta}(z_{t},t,c)},\end{split}$ (6)

### v2

@@ 3.1 Analysis on DDIM Inversion
$z_{t}=C_{t,1}*z_{t-1}+C_{t,2}*{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\epsilon_{\theta}(z_{t},t,c)},$ (1)
@@
The coupling of DDIM inversion with source prompt is rooted in the use of Eq. 2. In the ideal inversion (i.e., Eq. 1), the input should be $z_{t}$ and $t$, whereas previous methods have utilized $z_{t-1}$ and $t-1$ as the input (i.e., Eq. 2). However, $z_{t-1}$ is obtained by denoising $z_{t}$ conditioned with source prompt. Therefore, Eq. 2 actually introduces source prompt prior into the update. Taking the last inversion step as an example, the neural network should receive a latent $z_{T}$ a pure noise without any prior information as the input to obtain the ideal inversion code as $z_{T}=C_{T,1}*z_{T-1}+C_{T,2}*{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\epsilon_{\theta}(z_{T},T,c)}$, while most previous methods utilize the latent feature $z_{T-1}$, which is obtained through one-step denoising conditioned on the source prompt, as the input, i.e., $\hat{z}_{T}=C_{T,1}*z_{T-1}+C_{T,2}*{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\epsilon_{\theta}(z_{T-1},{T-1},c)}$. The coupling of the inverted noise code and source prompt is not limited to the final inversion step, as the multi-step generation nature of LDM accumulates the divergence between the inverted noise and the ideal noise. This eventually results in the inclusion of source prompt prior in the inverted noise, causing a notable deviation from the ideal noise, i.e., gap of noise $D_{Noi}$ as illustrated in Fig. 2(a). This deviation ultimately affects the editing stability and fidelity of the given real image.
@@ 3.2 Source Prompt Disentangled Inversion (SPDInv)
$x=f_{\theta}(x)\quad where\quad x=z_{t},f_{\theta}(x)=C_{t,2}*\epsilon_{\theta}(x,t,c)+C_{t,1}*z_{t-1}.$ (3)
@@ Appendix 0.A Reduction of Noise Gap by SPDInv
We conduct experiments to evaluate the noise gap reduction by the proposed SPDInv (please refer to Fig. 2 in the main paper and the analysis in the introduction section for details of noise gap $D_{noi}$). We first utilize the captions extracted from the COCO2017 evaluation dataset[26] to generate 100 images by Stable Diffusion V1.4. We then record the initial noises $z_{T}$ and all the latent features $z_{t},t\in[0,T)$, which can be regarded as the ground-truth features of the generated images (please refer to the generation path in Fig. 2 of our main paper). Since previous text-driven image editing methods mostly employ DDIM inversion to compute the inverted noise code, we thus compare DDIM inversion with our SPDInv on their inversion performance using the generated ground-truth features and conduct qualitative and quantitive experiments.
For qualitative evaluation, we utilize t-SNE to reduce the dimension of inverted noise codes and visualize them in Fig. 7. One can clearly see that the noise codes acquired through SPDInv exhibit a distribution more akin to Gaussian noise than DDIM inversion.
Figure 7: Visualization of inverted noise codes.
For quantitive evaluation, we first record all inverted latent features during the inversion process. Then for each inversion step $t$, we can calculate the mean square error (MSE) between the ground-truth and inverted latent features. Finally, the MSEs of all the 100 images are averaged as the $D_{noi}$, which is illustrated in Fig. 8. Compared with DDIM inversion, our SPDInv reduces the noise gap from 0.06 to 0.04 ($36\%\downarrow$) after 50 inversion steps. Based on the same setting, we further calculate the clip score between the Gaussian noise, DDIM inverted noise, SPDInv inverted noise and the source prompt. The results are shown in Tab. 4. We see that SPDInv can reduce the clip score by 10% compared with DDIM inversion, and it is closer to the clip score calculated by pure Gaussian noise. Owe to the reduced noise gap, our SPDInv can effectively reduce the effect of source prompt on the inverted noise code, and consequently reduce the editing artifacts.
@@
Table 4: Clip score between noise codes and source prompt.
Metrics Gaussian Noise DDIM SPDInv
Clip Score 12.66 14.55 13.18
@@
$\begin{split}z_{t-1}&=\frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}z_{t}+\sqrt{\alpha_{t-1}}\bigg{(}\sqrt{\frac{1}{\alpha_{t-1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1}\bigg{)}\epsilon_{\theta}(z_{t},t,c),\\ \frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}z_{t}&=z_{t-1}-\sqrt{\alpha_{t-1}}\bigg{(}\sqrt{\frac{1}{\alpha_{t-1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1}\bigg{)}\epsilon_{\theta}(z_{t},t,c),\\ \frac{\sqrt{\alpha_{t-1}}}{\sqrt{\alpha_{t}}}z_{t}&=z_{t-1}+\sqrt{\alpha_{t-1}}\bigg{(}\sqrt{\frac{1}{\alpha_{t}}-1}-\sqrt{\frac{1}{\alpha_{t-1}}-1}\bigg{)}\epsilon_{\theta}(z_{t},t,c),\\ z_{t}&=\frac{\sqrt{\alpha_{t}}}{\sqrt{\alpha_{t-1}}}z_{t-1}+\sqrt{\alpha_{t}}\bigg{(}\sqrt{\frac{1}{\alpha_{t}}-1}-\sqrt{\frac{1}{\alpha_{t-1}}-1}\bigg{)}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\epsilon_{\theta}(z_{t},t,c)},\end{split}$ (6)